Consider the following simplified financial statements for the Fire Corporation (assuming no income taxes): Income Statement Balance Sheet Sales $ 46,900 Assets $ 22,700 Debt $ 6,700 Costs 41,140 Equity 16,000 Net income $ 5,760 Total $ 22,700 Total $ 22,700 The company has predicted a sales increase of 18 percent. Assume Fire pays out half of net income in the form of a cash dividend. Costs and assets vary with sales, but debt and equity do not. Prepare the pro forma statements. (Round your answers to the nearest whole dollar amount.) Pro forma income statement Pro forma balance sheet Sales $ Assets $ Debt $ Costs Equity Net income $ Total $ Total $ Determine the external financing needed. (Negative amount should be indicated by a minus sign.) External financing needed $
Answer #1
  • Pro forma income statement
Sales    55,342
Costs    48,545
Net Income      6,797

Net income will be = 5760*1.18 = $ 6797

Retained Earning = 6797*50% = $ 3399

  • Pro forma balance sheet
Asset 26786 Debt 7387
Equity 19399
Total 26786 Total 26786
  • Determine the external financing needed

Total Fund Needed = 18%*22700 = $ 4086

Retained Earning = $ 3399

External Financing = 4086 - 3399

External Financing = $ 687